OpenStudy (anonymous):

Trey is estimating the weight of his mountain bike. The actual weight of his mountain bike is 16.7kg Trey's estimate is 15kg. Find the absolute error and the percent error of Trey's estimate. If necessary, round your answers to the nearest tenth.

OpenStudy (anonymous):

@HaileyD @Loser66 @k_lynn

OpenStudy (anonymous):

Can someone please help me ?

OpenStudy (anonymous):

the absolute error is 16.7-15 =1.7kg % error is error/original quantity 1.7/16.7=0.1017 *100 =10.17%